OpenStudy (kainui):

In base 6, the first 7 digits of any numbers x and y obey this... Why? ;)

OpenStudy (kainui):

\[\ (x1350213 + y4205344)^n = x^n1350213 + y^n4205344\]

OpenStudy (just_one_last_goodbye):

If I get the answer will you give me Owl bucks?

OpenStudy (kainui):

M A Y B E

OpenStudy (just_one_last_goodbye):

Liar e.e

OpenStudy (kainui):

I've gotta go to sleep, I'll be back to check up on this later. ;P

OpenStudy (ikram002p):

will try

ganeshie8 (ganeshie8):

Notice that in base-6, we have \(4205344 = 6^7-1350213 +1\) \[\begin{align}\ [x1350213 + y4205344]^n &= [6^7(x+y) +1350213 +4205344 ]^n\\~\\ &= [6^7(x+y) +6^7+1 ]^n\\~\\ &=6^7A + 1 \end{align}\] For the given number format, the right most 7 digit string is \(0000001\) no matter what \(x\) and \(y\) are.

OpenStudy (kainui):

Wait a sec that doesn't seem right, here's a counter example picked at random! \[(3*1350213 + 5*4205344)^2 = 0330124521\] all in base 6 still. I think you're onto some good ideas though.

OpenStudy (kainui):

Hmm whoops I put the answer in base 10 on the right there. I need a better way of going between base 6 and not. Something doesn't seem right overall, maybe I'm mistaken here.

OpenStudy (kainui):

Alright I'll give my solution since I worked it out after confusing myself and want to just get it out there before I forget. http://www.wolframalpha.com/input/?i=%5B(3%2B1*6%2B2*6%5E2%2B0*6%5E3%2B5*6%5E4%2B3*6%5E5%2B1*6%5E6)%5E2%5D+base+6 This shows that for the first several digits in base 6, \[(1350213)^2 = 1350213\] I'll go ahead and name it based on the digit it ends in, \(3\) as: \[i_3 = 1350213\] So it obeys:\[i_3^2=i_3\] Similarly the other number \[i_4=4205344\]obeys\[i_4^2=i_4\] Now the nail in the coffin is: \[i_3i_4=0\] So now it automatically will collapse this sucker down, all the cross terms go away!\[\ (x1350213 + y4205344)^n = x^n1350213 + y^n4205344\]

Parth (parthkohli):

Jesus man.

OpenStudy (kainui):

Cool right? If only we had numbers that infinitely extended to the left, then it would apply not just to the first 6 or so digits, it'd apply to all digits...!

OpenStudy (kainui):

Ok, so basically I cooked this question up after playing around with 6-adic numbers, these are approximations to some of those numbers. :P